OpenStudy (anonymous):

Which of the following is equal to C(10, 3)?

OpenStudy (anonymous):

C(3, 10) C(10, 7) P(10, 3)

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow C(10,3)={10! \over (10 - 3)! \times 3!} = {10! \over (10 - 7)! \times 3!} = C(10,7) }\)

Parth (parthkohli):

2nd option applies completely

OpenStudy (anonymous):

ok thanks